OpenStudy (anonymous):

-8 + log base 9 (m+1) =-8

OpenStudy (anonymous):

Look over your laws of logs. . .they are imperative to know for this unit. . . log(base 3) 45 - log(base 3) 9 = log_3 (45 / 9) = log_3 (5) and if solving. . .continue to. . . = log5 / log3 = 1.4645 Remember. . . log(base x)y + log(base x)z = log(base x)(y*z) log(base x)y - log(base x)z = log(base x)(y/z) log(base x)y = logy / logx log(base x)y^x/t = x/t(log(base x)y) The following site might be a great help to you. . . http://oakroadsystems.com/math/loglaws.h … log(base 2) 5 + log(base2) x - log(base 2) 10 = log(base 2)[(5x)/10] = log(base 2)0.5x log(base 2) 14 - log(base 2) 7 = log_2(14 / 7) = log_2(2) 4log(base b) m + (1/2)log (base b) n - 3log(base b) p = log_b(m)^4 + log_b(n)^(1/2) - log_b(p)^3 = log_b[(m^4*n^(1/2)) / (p^3)] = log_b [(m^4*(sqrt)n) / (p^3)]